OpenStudy (anonymous):

In a sample of 50 museum visitors one afternoon, 22 were female and 28 were male. If there were 800 museum visitors that afternoon, estimate the number of visitors who were female. a.220 b.352 c.448 d.700 GET A MEDAL

OpenStudy (twizttiez):

352 (22/50).800 = 22*16 = 352

OpenStudy (twizttiez):

Hope it helps ^-^

OpenStudy (anonymous):

it did , thank you @TwiztTiez